# OTTER: Improving Zero-Shot Classification via Optimal Transport

### v1

@@ 4.2 General Perturbation Sensitivity
In practical applications, calibration error could extend beyond noise in the elements of the cost matrix. A key source of error is label distribution estimation error. Hence, we address a more general setting, examining the impact of simultaneous perturbations in the label distribution and cost matrix of the transport plan. Our result applies techniques from the theory of linear programming perturbations.
@@
We adapt a theorem from Robinson [53] with our optimal transport problem notation.
@@
Let the primal linear programming problem be defined as in equation (1), and its dual problem be $max\{w^{T}g|w^{T}G\leq\text{vec}(C)^{T},w\geq 0\}$. Suppose perturbed cost matrix is $\hat{C}=C+\Delta_{C}$, the perturbed class distribution $\hat{\nu}=\nu+\Delta_{\nu}$, such that $\hat{g}=g+\Delta_{g}$ where
@@
$\begin{Vmatrix}\begin{pmatrix}\text{vec}(\pi)\\ w\end{pmatrix}-\begin{pmatrix}\text{vec}(\hat{\pi})\\ \hat{w}\end{pmatrix}\end{Vmatrix}_{2}\leq\kappa(\|\Delta_{\nu}\|_{2}+\|[\text{vec}(\Delta_{C})]_{+}\|_{2}+\|\text{vec}(C)^{T}\text{vec}(\hat{\pi})-g^{T}\hat{w}\|_{2}),$
where $1\leq p\leq\infty$ and $\kappa$ is a Hoffman constant that only relates to the original problem [28].
Ignoring the constant, the upper bound can be decomposed into three components,
@@
  • $\text{vec}(C)^{T}\text{vec}(\hat{\pi})-g^{T}\hat{w}$: the suboptimality of perturbed solutions $\hat{w},\text{vec}(\hat{\pi})$ in the original problem.
Theorem 4.3 implies that the deviation from perturbed solution to true solution is bounded by the magnitude of perturbations and suboptimality of the perturbed solution. From this result, we expect the prediction accuracy deteriorates linearly given the perturbations of the label distribution and the calibration.
@@ 5 Experiments
The primary objective of our experiments is to (1) validate that OTTER improves zero-shot model performance when given accurate label distribution estimates and (2) investigate its sensitivity to perturbations. In synthetic experiments (Section 5.1), we validate our theoretical claims—label shift invariance and sensitivity to perturbation in a fully controllable setting. In experiments on real datasets (Section 5.2), we confirm that OTTER can improve zero-shot classification significantly in a variety of settings (Section 5.2.1). In addition, we show that OTTER can be combined with label distribution estimation methods in the few-shot learning setting (Section 5.2.2). Finally, we show that OTTER especially benefits hierarchical classification (Section 5.2.3).
5.1 Synthetic Experiments
(a) Prediction accuracy changes with perturbed confidence score.
(b) Prediction accuracy changes with perturbed label distribution.
Claim Investigated. We hypothesize OTTER is invariant to label shift under the conditions in Theorem 4.2. We also investigate the sensitivity to perturbations of the cost matrix and the label distribution.
We separately investigate perturbed prediction score matrix and perturbed label distribution’s impact on the prediction accuracy. For perturbed prediction scores, we fix the label distribution to be the true one, and add noise $\delta\sim\mathcal{N}(0,\sigma^{2})$ of varying levels $\sigma$ to the predicted score $P_{\theta}(Y=1|X)$. For label distribution, we fix the prediction score to be true scores and add noise $\epsilon$: $\hat{\nu}=\nu^{t}+(\epsilon,-\epsilon)$. We use these perturbed variants to obtain perturbed solutions and compare with ground-truth solution.
Results. Figure 2 illustrates how accuracy changes with label shift when the predicted score is perturbed and when label distribution is perturbed. We observe that the naive classifier deteriorates as the total variation distance between source and target distributions increases. It indicates that naive classifier is sensitive to label shift. However, without perturbation, OTTER remains unaffected by the label distribution shift, which validates our invariance result in Section 4.
In the case of confidence score perturbation, both the naive classifier and OTTER have accuracy decreasing as perturbation level increases. For simplicity, we omitted the naive classifier’s performances under different levels of noise as adding zero-mean noise does not alter its accuracy significantly. We observe that OTTER has better performance than the naive method when significant label shift exists. Similarly, for label distribution perturbation, we observe as the noise level $\epsilon$ increases, OTTER’s accuracy downgrades—but still yields better performance when label shift is severe.
Our experimental results suggest simply applying predicted score for zero-shot classification leads to inaccurate predictions under label shift, while OTTER is robust to label shift when no perturbation existed. Perturbations in both predicted score and label distribution downgrades the predicted accuracy, as expected, but OTTER still yields better results than the naive baseline.
@@ 5.2 Real Data Experiments
Table 1: Accuracy (%) in zero-shot image classification (ViT-B/16). We use the true label distribution as the label distribution specification. The numbers in parenthesis of Prior Matching represent the standard deviation of 10 different samplings of the validation set.
5.2.1 Known label distribution
@@
Claim Investigated. We hypothesize that the model performance can improve significantly when the label distribution specification is exact.
@@
Setup and Procedure. We used 17 image classification datasets and 4 text classification datasets. We employed CLIP [48] for image zero-shot classification, and BERT [17]. A comprehensive list and details of experiments can be found in Appendix D.
@@
Baseline. We adopt Prior Matching (PM) [35] as a baseline. It optimizes score weighting parameters to align with the label distribution specification. A detailed explanation of Prior Matching is given in Appendix B. It is worth noting that the performance of Prior Matching is highly sensitive to hyperparameters such as temperature and learning rate. Optimal hyperparameters may vary across different datasets. We selected hyperparameters through grid search, by evaluating their performance on a validation set, consisting of 10 labeled examples per class. In contrast, we highlight that OTTER is tuning-free.
@@
Results. Table 1 shows the results for CLIP (ViT-B/16) and Table 2 shows the text classification results with BERT. Notably, OTTER demonstrates a 4.8% and 15.5% enhancement on average in image and text zero-shot classification, respectively. While Prior Matching shows competitive performance when accurately tuned, it is prone to failure. We found that hyperparameter tuning fails in the class-imbalanced datasets such as Caltech256, SUN397, ImageNet-r (Appendix D, Table 7). This suggests that the hyperparameter selection process necessitates a validation set label distribution similar to the target distribution—rendering unusable in zero-shot scenarios. Comprehensive results, including the sensitivity study on the label distribution specification error, are provided in Appendix D.3.
Zero-shot Prior Matching OTTER
Amazon review 74.0 58.8 (±$\pm 46.4$) 91.7
GenderBias 84.1 41.4 (±39.6$\pm$ 39.6)± 39.6 ) 91.9
CivilComments 48.4 57.2 (±$\pm 37.7$) 81.4
HateXplain 30.9 31.3 (±$\pm 3.3$) 34.3
Table 2: Accuracy (%) in zero-shot text classification with BERT. We use the true label distribution as the label distribution specification. For Prior Matching we report the standard deviation for 10 different samplings of the validation set.
5.2.2 Few-shot adaptation with label distribution estimation
Claim Investigated. We anticipate that OTTER can be used in few-shot learning when combined with label distribution estimation methods.
Setup and Procedure. We use the same datasets as the previous experiment. We consider a 10-shot learning setting: 10 labeled samples per class are given. Note that labeled samples have uniform label distribution, while the label distribution in the target distribution may not be uniform. This setting requires the use of label distribution estimation methods used in label shift adaptation [33, 6, 22]. We estimate the target label distribution with Black Box Shift Estimation (BBSE) [33]. BBSE estimates the target balance using confusion matrix, under the label shift assumption. For detailed explanation, refer to Appendix B.
Expected Results. We expect OTTER can improve zero-shot classification if the label distribution estimation error is sufficiently small. Also, we expect OTTER can improve linear probing, which is one of standard approaches for few-shot learning.
Results. We found OTTER improves zero-shot text classifications where the number of classes is small ($K=2$ or $3$), and improves linear probing in image classification with a large number of classes. Table 3 shows the text zero-shot classification results. While it shows a relatively high variance due to the small sample size (20 $\sim$ 30), the average accuracy significantly improves over zero-shot classification. Table 4 shows the linear probing experiment results. The result shows that OTTER can yield additional improvement upon linear probing. Full results that have both of zero-shot classification and linear probing are given in Appendix D.
Zero-shot Prior Matching OTTER
Amazon review 74.0 47.9 (±$\pm 45.5$) 89.9 (±$\pm 11.1$)
GenderBias 84.0 57.0 (±$\pm 36.5$) 87.8 (±$\pm 4.4$)
CivilComments 48.4 69.1 (±$\pm 33.7$) 55.81 (±$\pm 17.9$)
HateXplain 30.9 34.4 (±$\pm 6.0$) 35.2 (±$\pm 3.3$)
Table 3: Accuracy (%) in zero-shot text classification with BERT. We use the label distribution estimation by BBSE for Prior Matching and OTTER. We report the mean and the standard deviation of 10 different samplings of the labeled data.
LP LP + PM LP + OTTER
CIFAR10 90.2 (±$\pm 0.6$) 89.8 (±$\pm 1.1$) 90.0 (±$\pm 0.6$)
CIFAR100 58.3 (±$\pm 0.8$) 24.4 (±$\pm 12.2$) 60.5 (±$\pm 0.6$)
Caltech101 91.5 (±$\pm 0.8$) 87.5 (±$\pm 9.3$) 91.4 (±$\pm 0.9$)
Caltech256 84.5 (±$\pm 0.4$) 58.4 (±$\pm 23.5$) 85.4 (±$\pm 0.4$)
Country211 12.4 (±$\pm 0.4$) 9.2 (±$\pm 2.4$) 13.2 (±$\pm 0.3$)
DTD 58.6 (±$\pm 1.1$) 49.0 (±$\pm 9.3$) 59.3 (±$\pm 1.1$)
EUROSAT 74.6 (±$\pm 1.4$) 71.6 (±$\pm 3.5$) 75.9 (±$\pm 3.0$)
Flowers102 89.0 (±$\pm 0.0$) 87.8 (±$\pm 0.0$) 90.2 (±$\pm 0.1$)
Food101 79.1 (±$\pm 0.9$) 60.6 (±$\pm 17.1$) 79.8 (±$\pm 0.6$)
Oxford-IIIT-Pet 75.7 (±$\pm 1.3$) 72.0 (±$\pm 8.8$) 75.6 (±$\pm 1.7$)
Stanford-Cars 64.5 (±0.8$\pm$ 0.8)± 0.8 ) 65.4 (±$\pm 0.9$) 66.3 (±$\pm 0.8$)
STL10 97.7 (±$\pm 0.6$) 97.5 (±$\pm 0.9$) 97.6 (±$\pm 0.5$)
CUB 72.2 (±$\pm 0.1$) 63.3 (±$\pm 23.0$) 75.6 (±$\pm 0.4$)
ImageNet 56.8 (±$\pm 0.0$) 53.6 (±$\pm 4.1$) 59.8 (±$\pm 0.0$)
ImageNet-r 54.9 (±$\pm 0.8$) 47.6 (±$\pm 11.0$) 57.1 (±$\pm 0.8$)
ImageNet-Sketch 43.4 (±$\pm 0.0$) 37.9 (±1.4$\pm$ 1.4)± 1.4 ) 48.3 (±$\pm 0.0$)
Table 4: Accuracy (%) in linear probing image classification with ViT-B/16. We use the true label distribution as the label distribution specification. We report the meand and standard deviation of 10 different random samplings of the validation set. We excluded SUN397 since linear probing collapses.
Table 4 shows the experiment results. The result shows that OTTER can yield mild improvement over linear probing, even with the label distribution estimation errors. Figure 3 shows that accuracy improvement is consistent across the number of samples used for linear probing. More detailed analysis regarding estimation erorr and the number of samples is provided in Appendix D.4.
(a) BBSE label distribution estimation error
@@ 5.2.2 Few-shot adaptation with label distribution estimation
Figure 3: Ablation on the number of samples in few-shot learning.
5.2.3 Zero-shot prediction improvement with class hierarchy
Claim Investigated. We hypothesize incorporating class hierarchy information can enhance few-shot label distribution estimation and thus improve zero-shot predictions.
@@ 5.2.3 Zero-shot prediction improvement with class hierarchy
Table 5: Accuracy (%) with hierarchical OTTER (H-OTTER). We report the mean and the standard deviation of 10 different random samplings of the labeled data.
@@
6 Related work
Improving Zero-shot Classification at Inference Time. As zero-shot classification has gained popularity, several works have been developed to improve zero-shot classification at inference time. Chuang et al. [12], Adila et al. [1] use vector projection methods to remove spurious correlations at inference time. Menon and Vondrick [38], Novack et al. [42], An et al. [4] augment prompts with language models and combine their classification output to improve zero-shot performance. Roberts et al. [52] uses additional information of label space geometry to extend model pre-trained on the label subset to broader use-cases. While these works try to improve zero-shot classification at inference time in common, the main difference is that our method tackles the inherent class prior of zero-shot models.
Label Shift Adaptation. Label shift adaptation methods are designed to address the negative impacts arising from changes in the label distribution. These methods typically follow a two-step process [33, 6, 22]. The first step involves estimating the label distribution within the target dataset using labeled data from the source distribution and unlabeled data from the target distribution. Next, the prediction scores are reweighted using the estimated target label distribution and the source label distribution. However, the standard approach requires access to the labeled source distribution data, which is usually not possible in zero-shot classification scenarios. OTTER provides a solution decoupled from the source data distribution, overcoming this limitation.
Improving Zero-shot Classification using Prior. Several studies have explored leveraging prior information to enhance zero-shot classification, even in the absence of access to source distributions. In the context of prompt-based zero-shot models, prior matching [35] employs word prior distribution to alleviate word bias inherent in pretraining data. We adopted their adaptation method as a baseline. Similarly, Kahana et al. [29] develop adaptation layers trained using priors, aiming to maintain proximity to the original scores. However, both approaches entail training additional layers and necessitate hyperparameter tuning, which may pose challenges in the context of zero-shot predictions. In contrast, OTTER presents a straightforward and efficient adaptation method to new label distributions without the need for any hyperparameter tuning, backed by theoretical guarantees.
Leveraging Optimal Transport for Enhanced Pseudo-labeling and Classification. A number of studies have explored the enhancement of pseudo-labeling and classification tasks through optimal transport, using label distribution specifications, in a similar spirit to our work, but within different contexts. Tai et al. [60] uses optimal transport to allocate pseudo-labels to unlabeled datasets based on the label distribution specification in the semi-supervised setup. Wang et al. [64] deals with long-tail distribution in the partial-label learning setup based on optimal transport. Zhang et al. [70] uses partial optimal transport as a pseudo-labeling based approach for deep imbalanced clustering, progressively expanding the labeled sample proportion. Guo et al. [24] reweights the training dataset to match the label distribution specification using optimal transport. This work mainly deals with the class imbalance problem in the training step. Shi et al. [58] studies classification from a matching perspective, revealing the connection between softmax loss and inverse optimal transport and suggesting a new loss function to address long-tail distributions. Their analysis provides useful insights for OTTER — why cost matrix induced by pre-trained models can be useful in the inference step. Xian et al. [67] uses optimal transport as a postprocessing method to guarantee demographic parity. While sharing aspects of the approach, our work addresses class bias in zero-shot models. Peng et al. [46] used optimal transport to handle long-tail recognition with a learned cost matrix. Our study provides a theoretical basis for understanding their empirical results. Chang et al. [9] employs optimal transport to detect common and private classes between the source and the target domain, under the universal domain adaptation setting, where knowledge is transferred from source to target domain without any constraints on the label sets. In the context of zero-shot classification, there is no need to manage label space disparities between the source and target domains. Instead, the main concern of zero-shot classification is dealing with the distribution shift between the pretraining dataset and the downstream task. We tackle the label distribution shifts using optimal transport.
Class Imbalance. Class imbalance problems occur when the number of instances across different classes in a dataset is disproportionately distributed. This imbalance can severely bias the traininig process of a machine learning model, leading to poor generalization performance, especially for the minority classes. It has been extensively studied in the context of traditional machine learning [21]. Oversampling [11] and cost-sensitive learning [61] are well-known approaches to address class imbalance. Nonetheless, the inherent nature of class imbalance in pretraining datasets introduces a distinct set of challenges, especially when attempting to rectify such biases within the context of zero-shot classification scenarios.

### v2

@@ 4.2 General Perturbation Sensitivity
In practical applications, calibration error could extend beyond noise in the elements of the cost matrix. A key source of error is label distribution estimation error. Hence, we address a more general setting, examining the impact of simultaneous perturbations in the label distribution and cost matrix of the transport plan. Our result applies techniques from perturbation theory for linear programming .
@@
We adapt a theorem from robinson1973bounds with our optimal transport problem notation.
@@
Let the primal linear programming problem be defined as in equation (1), and its dual problem be $\max\{w^{T}g|w^{T}G\leq\text{vec}(C)^{T},w\geq 0\}$. Suppose perturbed cost matrix is $\hat{C}=C+\Delta_{C}$, the perturbed class distribution $\hat{\nu}=\nu+\Delta_{\nu}$, such that $\hat{g}=g+\Delta_{g}$ where
@@
$\|\pi-\hat{\pi}\|_{F}^{2}\leq\kappa^{2}\left(\|\Delta_{\nu}\|_{2}^{2}+\|[\text{vec}(\Delta_{C})]_{+}\|_{2}^{2}+\|\text{vec}(C)^{T}\text{vec}(\hat{\pi})-g^{T}\hat{w}\|_{2}^{2}\right)-\|w-\hat{w}\|_{2}^{2}$
, where $1\leq p\leq\infty$ and $\kappa$ is a Hoffman constant that only relates to the original problem [hoffman1952approximate].
Ignoring the constant and the subtraction part, the upper bound can be decomposed into three components,
@@
  • $\text{vec}(C)^{T}\text{vec}(\hat{\pi})-g^{T}\hat{w}$: the suboptimality of perturbed solution $\hat{w}.$
Theorem 4.3 implies that the deviation from perturbed solution to true solution is bounded by the magnitude of perturbations and suboptimality of the perturbed solution. From this result, we can expect prediction accuracy to deteriorate with perturbations in the label distribution and calibration.
[…]
We provide a theoretical result showing that R-OTTER can learn an optimal parameter by learning reweighting parameters from $\hat{y}_{\textsc{OTTER}}$ as pseudolabels, and produce identical predictions to OTTER  once the optimal parameter is obtained in the label shift setup.
Theorem 4.4.
Under the same assumptions as in Theorem 4.2, the parameter $r^{*}=P_{t}(Y)/P_{s}(Y)$ is optimal when learning with $y_{\textsc{OTTER}}$ as pseudolabels.
The proof is provided in Appendix C.6.
@@ 5 Experiments
[…]
@@ 5.1 Real Data Experiments
Table 1: Accuracy (%) in zero-shot image classification (ViT-B/16) and text classification (BERT). We use the true label distribution as the label distribution specification. The numbers in parenthesis of Prior Matching represent the standard deviation of 10 different samplings of the validation set. OTTER produces improvements nearly across-the-board, with an average lift 4.9% in image classification and 15.5% in text classification, outperforming a powerful baseline, prior matching in almost all cases.
@@
We hypothesize that the model performance can improve significantly when the label distribution specification is exact.
@@
Setup and Procedure. We used 17 image classification datasets and 4 text classification datasets. We employed CLIP [CLIP_base] for image zero-shot classification, and BERT [devlin2018bert]. A comprehensive list and details of experiments can be found in Appendix D.
@@
Baseline. We adopt Prior Matching (PM) [liusie2023mitigating] as a baseline. It optimizes score weighting parameters to align with the label distribution specification. A detailed explanation of Prior Matching is given in Appendix B. It is worth noting that the performance of Prior Matching is highly sensitive to hyperparameters such as temperature and learning rate. Optimal hyperparameters may vary across different datasets. We selected hyperparameters through grid search, by evaluating their performance on a validation set, consisting of 10 labeled examples per class. In contrast, we highlight that OTTER is tuning-free.
@@
Results. Table 1 shows the image classification results with CLIP (ViT-B/16) and the text classification results with BERT. Notably, OTTER demonstrates a 4.8% and 15.5% enhancement on average in image and text zero-shot classification, respectively. While Prior Matching shows competitive performance when accurately tuned, it often struggles. We found that hyperparameter tuning fails in the class-imbalanced datasets such as Caltech256, SUN397, ImageNet-r (Appendix D, Table 7). This suggests that the hyperparameter selection process necessitates a validation set label distribution similar to the target distribution—rendering it unusable in zero-shot scenarios. More details and additional experiment results — including the sensitivity study on the label distribution specification error, computation time, and combination with other prompting methods — are provided in Appendix D.3.
5.2 Synthetic Experiments
We hypothesize OTTER is invariant to label shift under the conditions in Theorem 4.2. We also investigate the sensitivity to perturbations of the cost matrix and the label distribution.
We separately investigate perturbed prediction score matrix and perturbed label distribution specification’s impact on the prediction accuracy. For perturbed prediction scores, we fix the label distribution to be the true one, and add noise $\delta\sim\mathcal{N}(0,\sigma^{2})$ of varying levels $\sigma$ to the predicted score $P_{\theta}(Y=1|X)$. For the perturbed label distribution specification, we fix the prediction score to be true scores and add noise $\epsilon$: $\hat{\nu}=\nu^{t}+(\epsilon,-\epsilon)$. We use these perturbed variants to obtain perturbed solutions and compare with ground-truth solution.
(a) Prediction accuracy changes with perturbed confidence score.
(b) Prediction accuracy changes with perturbed label distribution.
Results. Figure 2 illustrates how accuracy changes with label shift when the predicted score is perturbed and when label distribution specification is perturbed. We observe that the naive classifier deteriorates as the total variation distance between source and target distributions increases. It indicates that naive classifier is sensitive to label shift. However, without perturbation, OTTER remains unaffected by the label distribution shift, which validates our invariance result in Section 4.
In the case of confidence prediction perturbation, both the naive classifier and OTTER have accuracy decreasing as perturbation level increases. For simplicity, we omitted the naive classifier’s performances under different levels of noise as adding zero-mean noise does not alter its accuracy significantly. We observe that OTTER has better performance than the naive method when significant label shift exists. Similarly, for label distribution perturbation, we observe as the noise level $\epsilon$ increases, OTTER’s accuracy downgrades—but still yields better performance when label shift is severe.
Our experimental results suggest simply using prediction scores for zero-shot classification leads to inaccurate predictions under label shift, while OTTER is robust to label shift when no perturbations are present. Perturbations in both predicted score and label distribution specification downgrades the predicted accuracy, as expected, but OTTER still yields better results than the naive baseline.
Dataset ZS ZS BBSE+PM ZS BBSE+OT LP LP BBSE+PM LP BBSE+OT
CIFAR10 88.3 72.7 87.5 90.2 89.8 90.0
CIFAR100 63.8 3.2 59.1 58.3 24.4 60.5
Caltech101 79.8 32.5 80.7 91.5 87.5 91.4
Caltech256 79.8 6.0 80.3 84.5 58.4 85.4
Country211 19.8 1.5 15.9 12.4 9.2 13.2
DTD 39.0 3.2 31.2 58.6 49.0 59.3
EUROSAT 32.9 19.2 34.0 74.6 71.6 75.9
Flowers102 64.0 40.3 60.8 89.0 87.8 90.2
Food101 85.6 15.3 82.3 79.1 60.6 79.8
Oxford-IIIT-Pet 83.8 43.3 71.4 75.7 72.0 75.6
Stanford-Cars 55.7 2.3 51.7 64.5 65.4 66.3
STL10 98.0 97.4 96.9 97.7 97.5 97.6
SUN397 47.1 6.9 25.6 0.2 0.2 0.2
cub 46.0 3.3 45.5 72.2 63.3 75.6
ImageNet 60.2 0.8 57.7 56.8 53.6 59.8
ImageNet-r 68.9 1.7 63.3 54.9 47.6 57.1
ImageNet-Sketch 39.8 0.8 40.4 43.4 37.9 48.3
Amazon 74.0 47.9 89.1 71.3 66.9 71.3
CivilComments 48.3 69.1 55.8 53.8 45.5 53.8
Gender 84.0 57.0 87.8 78.0 71.2 78.5
HateXplain 30.4 34.4 35.2 32.8 32.7 32.3
Table 2: Accuracy (%) with OTTER combined with class balance estimation. ZS BBSE denotes BBSE label distribution estimation based on zero-shot prediction scores, and LP BBSE denotes BBSE label distribution estimation based on linear probing prediction scores. We report the mean of 10 different random samplings of the validation set. OTTER produces moderate improvements when comined with linear probing in image classification tasks. In text classification tasks, OTTER significantly improves accuracy, up to 15.1%, even with noisy label distribution estimation.
(a) BBSE estimation error
@@ 5.2 Synthetic Experiments
Figure 3: Ablation on the number of samples in few-shot learning. In (a), We can observe that BBSE estimation get more precise as the number of samples increases. Following this, OTTER gets better accuracy in (b). Additionally, OTTER consistently improves linear probing when combined.
5.3 Few-shot adaptation with label distribution estimation
We anticipate that OTTER can be used in few-shot learning when combined with label distribution estimation methods. We expect OTTER can improve zero-shot classification if the label distribution estimation error is sufficiently small. Also, we expect OTTER can improve linear probing, which is one of standard approaches for few-shot learning.
Setup and Procedure. We use the same datasets as the previous experiment. We consider a 10-shot learning setting: 10 labeled samples per class are given. Note that labeled samples have uniform label distribution, while the label distribution in the target distribution may not be uniform. This setting requires the use of label distribution estimation methods used in label shift adaptation [lipton2018detectingbbse, azizzadenesheli2019regularizedrlls, garg2020unifiedmlls]. We estimate the target label distribution with Black Box Shift Estimation (BBSE) [lipton2018detectingbbse]. BBSE estimates the target balance using confusion matrix, under the label shift assumption. For detailed explanation, refer to Appendix B.
Results. Table 2 shows the image and text zero-shot classification results with the label distribution estimation via BBSE and linear probing. The image classification results show that OTTER can yield mild improvement over linear probing, even with the label distribution estimation errors. Figure 3 shows that accuracy improvement is consistent across the number of samples used for linear probing. In text classification, we found OTTER improves zero-shot text classifications where the number of classes is small ($K=2$ or $3$). While it shows a relatively high variance due to the small sample size (20 $\sim$ 30), the average accuracy improves significantly over zero-shot classification. More detailed analysis regarding label distribution estimation error and the number of samples is provided in Appendix D.4.
@@ 5.3 Few-shot adaptation with label distribution estimation
Table 3: Accuracy (%) with hierarchical OTTER (H-OTTER). (H-OTTER) yields additional improvements over OTTER, up to 5.1%, using the hierarchy information of labels.
5.4 Zero-shot prediction improvement with class hierarchy
We hypothesize incorporating class hierarchy information can enhance few-shot label distribution estimation and thus improve zero-shot predictions.
@@ 5.4 Zero-shot prediction improvement with class hierarchy
[…]
Setup and Procedure.
[…]
Results.
[…]
Zero-shot OTTER R-OTTER Zero-shot OTTER R-OTTER
CIFAR10 88.3 91.7 88.4 Oxford-IIIT-Pet 83.8 88.8 85.7
CIFAR100 63.8 67.9 65.3 Stanford-Cars 55.7 59.7 51.0
Caltech101 79.8 88.7 88.2 STL10 98.0 98.6 98.1
Caltech256 79.8 87.0 79.6 SUN397 47.1 54.1 46.6
Country211 19.8 21.1 19.5 CUB 46.0 50.4 44.4
DTD 39.0 44.4 44.0 ImageNet 60.2 62.9 59.8
EUROSAT 32.9 42.0 39.3 ImageNet-r 68.8 72.4 68.5
Flowers102 64.0 70.8 69.7 ImageNet-Sketch 39.8 44.5 39.3
Food101 85.6 89.9 88.5
[…]
5.6 Mitigating selection bias in LLM multiple-choice questions
Selection bias is the tendency of LLMs to favor certain prefix tokens in multiple-choice questions [zheng2023selectionbias1, dalvidiscovering, choi2024selectionbias2, wei2024selectionbias3]. We demonstrate that OTTER can effectively mitigate selection bias by randomly shuffling the options and enforcing a uniform class balance in OTTER.
Setup and Procedure.
Our experimental setup follows zheng2023selectionbias1, utilizing the MMLU [hendrycks2020mmlu], ARC-Challenge [clark2018arcchallenge], and CommonsenseQA (CSQA) [talmor2019csqa] datasets. We test with llama-2(-chat)-7/13B [touvron2023llama] and vicuna-v1.3-7/13B [vicuna2023] models, treating the probabilities of each option token (A/B/C/D) as prediction probabilities. OTTER is applied under the assumption of a uniform distribution. We use 0-shot predictions and evaluate performance using accuracy and the standard deviation of recalls (RStd) as metrics.
Results.
ARC-Challenge CommonsenseQA (CSQA) MMLU
Naive OTTER Naive OTTER Naive OTTER
Model Acc.()($\uparrow$)( ↑ ) RStd ()($\downarrow$)( ↓ ) Acc. RStd Acc. RStd Acc. RStd Acc. RStd Acc. RStd
Llama-2-7b 36.0 27.4 45.5 1.7 31.9 28.4 42.7 3.8 36.1 22.5 40.0 0.9
Llama-2-13b 62.9 6.0 62.8 1.5 57.0 10.2 58.1 2.0 51.1 6.9 51.8 1.3
[…]
Llama-2-13b-chat 64.4 13.7 66.2 2.5 64.0 9.8 66.4 1.8 52.3 13.9 53.8 1.0
[…]
vicuna-13b 62.9 8.3 63.3 2.4 63.4 12.9 64.5 3.1 50.5 9.9 50.7 1.1
Table 5: Mitigation of selection bias via OTTER.
Table 5 presents the experimental results. OTTER significantly reduces selection bias, enhancing accuracy by up to 10.6% and lowering RStd by as much as 25.7%.